# Deeper Insights into Deep Graph Convolutional Networks: Stability and Generalization

### v1

@@ 4.2 Generalization Gap
This section presents the main results of this paper. For convenience, the notations used in the result are summarized in Table 1. Under the assumptions made in Section 4.1, the bound on the generalization gap of deep GCNs is provided in the following theorem.
@@
A fundamental correlation between the generalization gap and the parameters governing deep GCNs is induced by Theorem 1. This correlation implies that the uniform stability of deep GCNs, trained using the SGD algorithm, exhibits an increase with the number of samples when the upper bound approaches zero as the sample size $m$ tends to infinity. Specifically, it is observed that if the value of $C_{g}$ (presenting the largest absolute eigenvalue of a symmetry $g(\mathbf{L})$ or the maximum singular value of an asymmetry $g(\mathbf{L})$) remains unaffected by the size $N$, a generalization gap decaying at the order of $O(1/\sqrt{m})$ is obtained. To compare with the result in [36], let us discuss at length the role of $g(\mathbf{L})$ and the hidden layer number $K$ on the generalization gap.
@@
$\epsilon_{gen}(\mathcal{A}_{\mathcal{S}})\leq\frac{1}{\sqrt{m}}\bigg{(}O\Big{(}C_{g}^{2T(K+1)}\Big{)}+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\bigg{)}.$ (9)
When $K=0$, the GCN model (1) degenerates into the single-layer GCN model considered in [36]. At this point, according to (9), we have
$\epsilon_{gen}(\mathcal{A}_{\mathcal{S}})\leq\frac{1}{\sqrt{m}}\bigg{(}O\Big{(}C_{g}^{2T}\Big{)}+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\bigg{)},$ (10)
which is the same as the result of [36].
@@
    Normalized vs. Unnormalized Graph Filters: We examine the three most commonly utilized filters: 1) $g_{1}(\mathbf{L})=\mathbf{A}+\mathbf{I}$, 2) $g_{2}(\mathbf{L})=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}+\mathbf{I}$, and 3) $g_{3}(\mathbf{L})=\mathbf{D}^{-1}\mathbf{A}+\mathbf{I}$. For the unnormalized filter $g_{1}$, its maximum absolute eigenvalue is bounded by $O(N)$. Consequently, as the value of $m$ approaches the magnitude to $N$, the upper bound indicated by (9) tends towards $O(N^{p})$ for some $p>0$, leading to an impractical upper bound when $N$ become infinitely large. On the contrary, for two normalized filters $g_{2}$ and $g_{3}$, their largest absolute eigenvalues are bounded and independent of graph size $N$. Therefore, both filters yield a diminishing generalization gap at a rate of $O(\frac{1}{\sqrt{m}})$ as $m$ goes to infinity. This discovery underscores the superior performance of normalized filters over unnormalized counterparts in deep GCNs. This observation is consistent with the findings in [36, 37].
@@
    The Role of Parameter $K$: It is evident that, when the values of $C_{g}$ and $T$ are fixed, the upper bound (9) exhibits an exponential dependence on parameter $K$. This observation implies that a larger value $K$ leads to an increase in the upper bound of the generalization gap, thereby offering valuable insights for the architectural design of deep GCNs. This finding diverges from the ones presented in [36, 37], as these studies do not account for generic deep GCNs and overlook the significance of the parameter $K$.
@@
Ref. Architecture
Estimated Upper Bound of the Generalization Gap
[36] shallow $\frac{1}{\sqrt{m}}\bigg{(}O\Big{(}(1+\eta\upsilon_{\ell}\upsilon_{\sigma}C_{g}^{2})^{T}\Big{)}+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\bigg{)}$
[37]
shallow
$\frac{1}{\sqrt{m}}\bigg{(}O\Big{(}\eta\alpha_{\ell}\alpha_{\sigma}c_{2,T}\sum\limits_{t=0}^{T-1}c_{6,t}\prod\limits_{s=t+1}^{T-1}(1+\eta c_{5,s})\Big{)}+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\bigg{)}$
[42] shallow $\frac{1}{\sqrt{m}}\Bigg{\{}O\bigg{(}C^{2}_{g}\eta C_{p,\lambda}\sum\limits_{t=1}^{T}(C_{p,\lambda}(1+(\alpha_{\sigma}^{2}+\alpha_{\ell})\eta C_{g}^{2}))^{t-1}\bigg{)}+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\Bigg{\}}$
@@
    Note: $\delta\in(0,1)$, $c_{2,t}$, $c_{6,t}$ and $c_{5,t}$ ($t=0,1,\dots,T$) represent some specific parameters defined in [37]; $C_{p,\lambda}=\frac{28}{p(p-1)\lambda_{t}}(B/\lambda)^{(3-p)/p}$ where $1<p\leq 2$, $\lambda>0$ is the regularization parameter and $\lambda_{t}>0$ is another regularization parameter dependent on $\lambda$ and $t$, as detailed in [42]. For information on other parameters, refer to Table 1.
Table 2 offers a concise summary of various upper bounds on the generalization gap, derived through the application of uniform stability. From Table 2, we can see that all the works derive a generalization gap decaying at the order of $O(1/\sqrt{m})$. However, compared to the other three works which only consider shallow GCNs, our work explores the case of deep GCNs. We should point out that the generalization of single-layer GCNs into deep GCNs is not trivial. To derive the results for deep GCNs, we tackle two significant challenges that arise specifically in the context of deep GCNs, which are unique to deep GCNs and are non-existent in single-layer models. The first challenge is the derivation of the gradient of the final output with respect to the learnable parameters across multiple layers, which requires determining how the gradient of the overall error of a GCN is shared among neurons in different hidden layers. In particular, in Appendix A.1, we provide a recursive formula to compute the related gradients. The second challenge is the evaluation of gradient variations between GCNs trained on different datasets. In the single layer case, since the input feature is the same, the variation of the related gradient is only dependent on the variations of learnable parameters. While, in the case of deep GCNs, the variation of the related gradients is also dependent on the variations of the gradients of the final output with respect to the hidden layer outputs. Please see Lemma 7 and its proof for details.
@@ 4.3 Stability Upper Bound
Let us consider the deep GCNs defined by equation (1). These networks are trained on a dataset $\mathcal{S}$ using the SGD algorithm for a total of $T$ iterations and denoted as $\mathcal{A}_{\mathcal{S}}$. Assume that Assumptions 1, 2 and 3 stated in Section 4.1 are satisfied. Then, $\mathcal{A}_{\mathcal{S}}$ is $\mu_{m}$-uniformly stable, with $\mu_{m}$ satisfying the following condition:
@@
Consider $\mathcal{A}_{\mathcal{S}}$, a set of deepGCNs defined by (1), trained on the dataset $\mathcal{S}$ using SGD for $T$ iterations. Let $\theta_{t}=\{\mathbf{W}^{(1)}_{t},\dots,\mathbf{W}^{(K)}_{t},\mathbf{w}_{t}\}$ and $\theta^{\prime}_{t}=\{\mathbf{W}^{(1)^{\prime}}_{t},\dots,\mathbf{W}^{(K)^{\prime}}_{t},\mathbf{w}^{\prime}_{t}\}$ denote the parameters of two GCNs trained on $\mathcal{S}$ and $\mathcal{S}^{i}$ after $t$ iterations, respectively. We set $\Delta\mathbf{w}_{t}=\mathbf{w}_{t}-\mathbf{w}_{t}^{\prime}$ and $\Delta\mathbf{W}^{(k)}_{t}=\mathbf{W}^{(k)}_{t}-\mathbf{W}^{(k)^{\prime}}_{t}$ to be the perturbation of learning parameters and define
$\|\Delta\theta_{t}\|_{*}=\|\Delta\mathbf{w}_{t}\|_{2}+\sum_{k=1}^{K}\|\Delta\mathbf{W}_{t}^{(k)}\|_{2}.$ (12)
@@
We provide the proof of Lemma 2 in A.2.
@@
$\triangle\mathbf{W}_{t}^{(k)}=\triangle\mathbf{W}^{(k)}_{t-1}-\eta\Big{(}\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}_{t}|\theta_{t-1}),y_{t})-\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}^{\prime}_{t}|\theta_{t-1}^{\prime}),y^{\prime}_{t})\Big{)}.$
Then, we provide two Lemmas to bound
$\nabla_{\mathbf{w}}\ell(f(\mathbf{x}_{t}|\theta_{t-1}),y_{t})-\nabla_{\mathbf{w}}\ell(f(\mathbf{x}^{\prime}_{t}|\theta^{\prime}_{t-1}),y^{\prime}_{t})$
@@
$\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}_{t}|\theta_{t-1}),y_{t})-\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}^{\prime}_{t}|\theta^{\prime}_{t-1}),y^{\prime}_{t})$
in two cases of $(\mathbf{x}_{t},y_{t})=(\mathbf{x}^{\prime}_{t},y^{\prime}_{t})$ and $(\mathbf{x}_{t},y_{t})\neq(\mathbf{x}^{\prime}_{t},y^{\prime}_{t})$, as shown in Lemma 3 and Lemma 4.
@@
where $\kappa_{1}$ and $\rho_{k}$ are defined by (7) and (30), respectively.
@@
The proofs of Lemma 3 and Lemma 4 are given in A.3. Using Lemma 3 and Lemma 4, we now provide a bound for $\mathbb{E}_{\mathcal{A}}\big{[}\|\triangle\theta_{T}\|_{*}\big{]}$.
@@
The proof of Lemma 2 is provided in A.4. Combining (14) and Theorem 3, we obtain that the uniform stability $\mu_{m}$ of $\mathcal{A}_{\mathcal{S}}$ has a bound as
@@ 5 Experiments
In this section, we conduct some empirical studies using three benchmark datasets commonly utilized for the node classification task, namely Cora, Citeseer, and Pubmed [53, 54]. Table 3 summarizes the basic statistics of these datasets. In our experiments, we follow the standard transductive learning problem formulation and the training/test setting used in [55]. To rigorously test our theoretical insights, our experiments aim to answer the following key questions:
@@
To address each question, we empirically estimate the generalization gap by calculating the absolute difference in loss between training and test samples. We adopt the official TensorFlow implementation [^1] https://github.com/tkipf/gcn for GCN [55] and the Adam optimizer with default settings. The number of iterations is fixed to $T=200$ for all the simulations.
Table 3: Statistics of the three benchmark datasets.
Cora Citeseer Pubmed
# Nodes $2,708$ $3,327$ $19,717$
# Edges $5,429$ $4,732$ $44,338$
# Features $1,433$ $3,703$ $500$
# Classes $7$ $6$ $3$
Label Rate $0.052$ $0.036$ $0.003$
Results and Discussion for Q1. We analyze two types of graph filters in our study: 1) the normalized graph filter, defined as $g(\mathbf{L})=\mathbf{\tilde{D}}^{-1/2}\mathbf{\tilde{A}}\mathbf{\tilde{D}}^{-1/2}$ with $\mathbf{\tilde{A}}=\mathbf{A}+\mathbf{I}$ and $\mathbf{\tilde{D}}_{ii}=\sum_{j}\mathbf{\tilde{A}}_{ij}$ (which was first employed in the vanilla GCN [55] and has subsequently become widely used in follow-up works on GCNs), and 2) the random walk filter, $g(\mathbf{L})=\mathbf{D}^{-1}\mathbf{A}+\mathbf{I}$. To fit our theoretical finding, we compare the performance of two 5-layer GCN models (with width $d=32$ for each layer), each employing one of these filters. Table 4 presents the numerical records of $R_{emp}(\mathcal{A}_{\mathcal{S}})$, $R(\mathcal{A}_{\mathcal{S}})$, $\epsilon_{gen}(\mathcal{A}_{\mathcal{S}})$, $C_{g}$ for both filters. The results indicate clearly that the 5-layer GCN with the normalized graph filter exhibits a smaller generalization gap compared to the one with the random walk filter. Furthermore, Figure 1 illustrates the performance of each filter across different datasets over iterations, demonstrating the superior performance of the normalized graph filter. Overall, the empirical findings in Table 4 and Figure 1 align well with our theoretical finding regarding the impact of $C_{g}$ on the generalization gap.
@@
Results and Discussion for Q2. In this experimental study, we try different settings of $K$, i.e., the number of hidden layers. Specifically, for $K=\{1,2,3,4,5\}$, we compare the performance of two $K$-layer GCNs (with width $d=32$ for each layer): one employing the normalized graph filter $g(\mathbf{L})=\mathbf{\tilde{D}}^{-1/2}\mathbf{\tilde{A}}\mathbf{\tilde{D}}^{-1/2}$, and one using the random walk filter $g(\mathbf{L})=\mathbf{D}^{-1}\mathbf{A}+\mathbf{I}$. Figure 2 shows the performance comparison results for each $K$. It demonstrates clearly that, consistent with the aforementioned results for Q1, GCN with a normalized graph filter (with smaller $C_{g}$) consistently exhibits smaller generalization gaps compared to those with the random walk filter. Also, it is observed that the generalization gap becomes larger as $K$ increases, further validating our theoretical assertions regarding the influence of $K$ on the model’s generalization gap.

### v2

@@ 4.2 Generalization Gap
This section presents the main results of this paper. Under the assumptions made in Section 4.1, the bound on the generalization gap of deep GCNs is provided in the following theorem.
@@
A fundamental correlation between the generalization gap and the parameters governing deep GCNs is induced by Theorem 1. This correlation implies that the uniform stability of deep GCNs, trained using the SGD algorithm, exhibits an increase with the number of samples when the upper bound approaches zero as the sample size $m$ tends to infinity. Specifically, it is observed that if the value of $C_{g}$ (presenting the largest absolute eigenvalue of a symmetric $g(\mathbf{L})$ or the maximum singular value of an asymmetric $g(\mathbf{L})$) remains unaffected by the size $N$, a generalization gap decaying at the order of $O(1/\sqrt{m})$ is obtained. To compare with the result in 2019Stable_GCN , let us discuss at length the role of $g(\mathbf{L})$ and the hidden layer number $K$ on the generalization gap.
@@
$\epsilon_{gen}(\mathcal{A}_{\mathcal{S}})\leq\frac{1}{\sqrt{m}}\left(O\Big(C_{g}^{2T(K+1)}\Big)+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\right).$ (9)
When $K=0$, the GCN model (1) degenerates into the single-layer GCN model considered in 2019Stable_GCN . At this point, according to (9), we have
$\epsilon_{gen}(\mathcal{A}_{\mathcal{S}})\leq\frac{1}{\sqrt{m}}\left(O\Big(C_{g}^{2T}\Big)+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\right),$ (10)
which is the same as the result of 2019Stable_GCN .
@@
    Normalized vs. Unnormalized Graph Filters: We examine the three most commonly utilized filters: 1) $g_{1}(\mathbf{L})=\mathbf{A}+\mathbf{I}$, 2) $g_{2}(\mathbf{L})=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}+\mathbf{I}$, and 3) $g_{3}(\mathbf{L})=\mathbf{D}^{-1}\mathbf{A}+\mathbf{I}$. For the unnormalized filter $g_{1}$, its maximum absolute eigenvalue is bounded by $O(N)$. Consequently, as the value of $m$ approaches the magnitude to $N$, the upper bound indicated by (9) tends towards $O(N^{p})$ for some $p>0$, leading to an impractical upper bound when $N$ become infinitely large. On the contrary, for two normalized filters $g_{2}$ and $g_{3}$, their largest absolute eigenvalues are bounded and independent of graph size $N$. Therefore, both filters yield a diminishing generalization gap at a rate of $O(\frac{1}{\sqrt{m}})$ as $m$ goes to infinity. This discovery underscores the superior performance of normalized filters over unnormalized counterparts in deep GCNs. This observation is consistent with the findings in 2019Stable_GCN , 2023Stable_GCN .
@@
    Low-pass vs. High-pass Graph Filters: Our theoretical results are not restricted to the choice of $g(\mathbf{L})$ as either a low-pass or a high-pass filter. To illustrate, consider two exponential filters with symmetric $\mathbf{L}$: i) a low-pass filter $g_{\mathrm{low}}(\lambda)=e^{-b\lambda^{2}}$ and ii) a high-pass filter $g_{\mathrm{high}}(\lambda)=1-e^{-a\lambda^{2}}$, where $a,b>0$. In this setting, it is straightforward to verify that
    $\|g_{\textrm{high}}(\mathbf{L})\|_{2}<\|g_{\textrm{low}}(\mathbf{L})\|_{2}=1.$
    Consequently, both filters lead to a vanishing generalization gap at the rate of $O\left(\tfrac{1}{\sqrt{m}}\right)$ as $m\to\infty$.
  • 3.
    The Role of Parameter $K$: It is evident that, when the values of $C_{g}$ and $T$ are fixed, the upper bound (9) exhibits an exponential dependence on parameter $K$. This observation implies that a larger value $K$ leads to an increase in the upper bound of the generalization gap, thereby offering valuable insights for the architectural design of deep GCNs. This finding diverges from the ones presented in 2019Stable_GCN , 2023Stable_GCN , as these studies do not account for generic deep GCNs and overlook the significance of the parameter $K$.
@@
   Reference    Model Architecture    Estimated Upper Bound of the Generalization Gap
   2019Stable_GCN    shallow    $\frac{1}{\sqrt{m}}\bigg(O\Big((1+\eta\upsilon_{\ell}\upsilon_{\sigma}C_{g}^{2})^{T}\Big)+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\bigg)$
   2023Stable_GCN    shallow    $\frac{1}{\sqrt{m}}\bigg(O\Big(\eta\alpha_{\ell}\alpha_{\sigma}c_{2,T}\sum\limits_{t=0}^{T-1}c_{6,t}\prod\limits_{s=t+1}^{T-1}(1+\eta c_{5,s})\Big)+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\bigg)$
   2023Stable_Lp    shallow    $\frac{1}{\sqrt{m}}\Bigg\{O\bigg(C^{2}_{g}\eta C_{p,\lambda}\sum\limits_{t=1}^{T}(C_{p,\lambda}(1+(\alpha_{\sigma}^{2}+\alpha_{\ell})\eta C_{g}^{2}))^{t-1}\bigg)+M\sqrt{\frac{\log\frac{1}{\delta}}{2}}\Bigg\}$
@@
    Note: $m$ is the number of samples in the trained dataset; $M$ is the upper bound of loss function $\ell(\cdot,\cdot)$; $\eta>0$ is the learning rate; $\delta\in(0,1)$; $T$ is the number of iterations for training $\mathcal{A}_{\mathcal{S}}$ using SGD; $C_{g}$ represents the 2-norm of filter $g(\mathbf{L})$; $\alpha_{\sigma}$, $\upsilon_{\sigma}$ are two parameters w.r.t the continuity of activation function $\sigma(\cdot)$; $\alpha_{\ell}$, $\upsilon_{\ell}$ are two parameters w.r.t the continuity of the loss function $\ell(\cdot,\cdot)$. $c_{2,t},c_{6,t},c_{5,t}>0$ ($t=0,1,\dots,T$) represent some specific parameters defined in 2023Stable_GCN . $C_{p,\lambda}=\frac{28}{p(p-1)\lambda_{t}}(B_{\ell}/\lambda)^{(3-p)/p}$, where $B_{\ell}>0$ is a parameter related to loss function $\ell(\cdot,\cdot)$, $1<p\leq 2$, $\lambda>0$ is the regularization parameter and $\lambda_{t}>0$ is another regularization parameter dependent on $\lambda$ and $t$, as detailed in 2023Stable_Lp . $K$ is number of hidden layers of the considered deep GCNs; $\kappa_{1}$ and $\kappa_{2}$ are two parameters as defined in (7) and (8).
Table 2 offers a concise summary of various upper bounds on the generalization gap, derived through the application of uniform stability. From Table 2, we can see that all the works derive a generalization gap decaying at the order of $O(1/\sqrt{m})$. However, compared to the other three works which only consider shallow GCNs, our work explores the case of deep GCNs. We should point out that the generalization of single-layer GCNs into deep GCNs is not trivial. To derive the results for deep GCNs, we tackle two significant challenges that arise specifically in the context of deep GCNs, which are unique to deep GCNs and are non-existent in single-layer models. The first challenge is the derivation of the gradient of the final output with respect to the learnable parameters across multiple layers, which requires determining how the gradient of the overall error of a GCN is shared among neurons in different hidden layers. In particular, in Appendix A, we provide a recursive formula to compute the related gradients. The second challenge is the evaluation of gradient variations between GCNs trained on different datasets. In the single layer case, since the input feature is the same, the variation of the related gradient is only dependent on the variations of learnable parameters. While, in the case of deep GCNs, the variation of the related gradients is also dependent on the variations of the gradients of the final output with respect to the hidden layer outputs. Please see Lemma 7 and its proof for details (see Appendix C).
@@ 4.3 Stability Upper Bound
Consider the deep GCNs defined by equation (1), which are trained on a dataset $\mathcal{S}$ using the SGD algorithm for a total of $T$ iterations and denoted as $\mathcal{A}_{\mathcal{S}}$. Assume that Assumptions 1, 2 and 3 stated in Section 4.1 are satisfied. Then, $\mathcal{A}_{\mathcal{S}}$ is $\mu_{m}$-uniformly stable, with $\mu_{m}$ satisfying the following condition:
@@
Consider $\mathcal{A}_{\mathcal{S}}$, a set of deepGCNs defined by (1), trained on the dataset $\mathcal{S}$ using SGD for $T$ iterations. Let $\theta_{t}=\{\mathbf{W}^{(1)}_{t},\dots,\mathbf{W}^{(K)}_{t},\mathbf{w}_{t}\}$ and $\theta^{\prime}_{t}=\{\mathbf{W}^{(1)^{\prime}}_{t},\dots,\mathbf{W}^{(K)^{\prime}}_{t},\mathbf{w}^{\prime}_{t}\}$ (with $\theta_{0}=\theta^{\prime}_{0}$) denote the parameters of two GCNs trained on $\mathcal{S}$ and $\mathcal{S}^{i}$ after $t$ iterations, respectively. We set $\triangle\mathbf{w}_{t}=\mathbf{w}_{t}-\mathbf{w}_{t}^{\prime}$ and $\triangle\mathbf{W}^{(k)}_{t}=\mathbf{W}^{(k)}_{t}-\mathbf{W}^{(k)^{\prime}}_{t}$ to be the perturbation of learning parameters and define
$\|\triangle\theta_{t}\|_{*}=\|\triangle\mathbf{w}_{t}\|_{2}+\sum_{k=1}^{K}\|\triangle\mathbf{W}_{t}^{(k)}\|_{2}.$ (12)
@@
We provide the proof of Lemma 2 in Appendix B.
@@
$\triangle\mathbf{W}_{t}^{(k)}=\triangle\mathbf{W}^{(k)}_{t-1}-\eta\Big(\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}_{t}|\theta_{t-1}),y_{t})-\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}^{\prime}_{t}|\theta_{t-1}^{\prime}),y^{\prime}_{t})\Big),$
with $\|\triangle\theta_{0}\|_{*}=0$.
So, we need to bound
$\nabla_{\mathbf{w}}\ell(f(\mathbf{x}_{t}|\theta_{t-1}),y_{t})-\nabla_{\mathbf{w}}\ell(f(\mathbf{x}^{\prime}_{t}|\theta_{t-1}^{\prime}),y^{\prime}_{t})$
@@
$\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}_{t}|\theta_{t-1}),y_{t})-\nabla_{\mathbf{W}^{(k)}}\ell(f(\mathbf{x}^{\prime}_{t}|\theta_{t-1}^{\prime}),y^{\prime}_{t})$
to obtain a bound of $\|\triangle\theta_{t}\|_{*}$. There are two scenarios to consider: i) At step $t$, SGD picks a sample $\mathbf{z}_{t}=(\mathbf{x}_{t},\mathbf{y}_{t})$ which is identical in $\mathcal{S}$ and $\mathcal{S}^{i}$, and occurs with probability (m1)/m(m-1)/m; and ii) At step $t$, SGD picks the only samples that $\mathcal{S}$ and $\mathcal{S}^{i}$ differ, $\mathbf{z}_{t}=(\mathbf{x}_{t},\mathbf{y}_{t})$ and $\mathbf{z}^{\prime}_{t}=(\mathbf{x}^{\prime}_{t},\mathbf{y}^{\prime}_{t})$ which occurs with probability $1/m$. We provide the results in the following Lemma 3 and Lemma 4.
@@
where $\kappa_{1}$ and $\rho_{k}$ are defined by (7) and (A.12).
@@
The proofs of Lemma 3 and Lemma 4 are given in Appendix C. We now provide a bound for $\mathbb{E}_{\mathcal{A}}\big[\|\triangle\theta_{T}\|_{*}\big]$.
@@
The proof of Theorem 3, using Lemma 3 and Lemma 4, is provided in Appendix D. Combining (14) and Theorem 3, we obtain that the uniform stability $\mu_{m}$ of $\mathcal{A}_{\mathcal{S}}$ has a bound as
@@ 5 Experiments
In this section, we conduct some empirical studies using three benchmark datasets commonly utilized for the node classification task, namely Cora, Citeseer, and Pubmed sen2008collective , yang2016revisiting . Table 3 summarizes the basic statistics of these datasets.
Table 3: Statistics of the three benchmark datasets.
Cora Citeseer Pubmed
# Nodes $2,708$ $3,327$ $19,717$
# Edges $5,429$ $4,732$ $44,338$
# Features $1,433$ $3,703$ $500$
# Classes 77 66 33
Label Rate $0.052$ $0.036$ $0.003$
In our experiments, we follow the standard transductive learning problem formulation and the training/test setting used in gcn . To rigorously test our theoretical insights, our experiments aim to answer the following key questions:
@@
To address each question, we empirically estimate the generalization gap by calculating the absolute difference in loss between training and test samples. We adopt the official TensorFlow implementation (https://github.com/tkipf/gcn) for GCN gcn and the Adam optimizer with default settings. The number of iterations is fixed to $T=200$ for all the simulations.
Results and Discussion for Q1. We analyze two types of graph filters in our study: 1) the normalized graph filter, defined as $g(\mathbf{L})=\mathbf{\tilde{D}}^{-1/2}\mathbf{\tilde{A}}\mathbf{\tilde{D}}^{-1/2}$ with $\mathbf{\tilde{A}}=\mathbf{A}+\mathbf{I}$ and $\mathbf{\tilde{D}}_{ii}=\sum_{j}\mathbf{\tilde{A}}_{ij}$ (which was first employed in the vanilla GCN gcn and has subsequently become widely used in follow-up works on GCNs), and 2) the random walk filter, $g(\mathbf{L})=\mathbf{D}^{-1}\mathbf{A}+\mathbf{I}$. To fit our theoretical finding, we compare the performance of two 5-layer GCN models (with width $d=32$ for each layer), each employing one of these filters. Table 4 presents the numerical records of $R_{emp}(\mathcal{A}_{\mathcal{S}})$, $R(\mathcal{A}_{\mathcal{S}})$, $\epsilon_{gen}(\mathcal{A}_{\mathcal{S}})$, $C_{g}$ for both filters. The results indicate clearly that the 5-layer GCN with the normalized graph filter exhibits a smaller generalization gap compared to the one with the random walk filter. Furthermore, Fig. 1 illustrates the performance of each filter across different datasets over iterations, demonstrating the superior performance of the normalized graph filter. Overall, the empirical findings in Table 4 and Fig. 1 align well with our theoretical finding regarding the impact of $C_{g}$ on the generalization gap.
@@
Results and Discussion for Q2. In this experimental study, we try different settings of $K$, i.e., the number of hidden layers. Specifically, for $K=\{1,2,3,4,5\}$, we compare the performance of two $K$-layer GCNs (with width $d=32$ for each layer): one employing the normalized graph filter $g(\mathbf{L})=\mathbf{\tilde{D}}^{-1/2}\mathbf{\tilde{A}}\mathbf{\tilde{D}}^{-1/2}$, and one using the random walk filter $g(\mathbf{L})=\mathbf{D}^{-1}\mathbf{A}+\mathbf{I}$. Fig. 2 shows the performance comparison results for each $K$. It demonstrates clearly that, consistent with the aforementioned results for Q1, GCN with a normalized graph filter (with smaller $C_{g}$) consistently exhibits smaller generalization gaps compared to those with the random walk filter. Also, it is observed that the generalization gap becomes larger as $K$ increases, further validating our theoretical assertions regarding the influence of $K$ on the model’s generalization gap.